Is there a way or maybe an equation to tell what an imaginary number is going to be?

For example: i to the 14th power would be...

Answer by scott8148(6628)

i^0=1
i^1=i
i^2=-1
i^3=-i
i^4=1
i^5=i

i^14=-1 ... see the pattern? ... it repeats indefinitely